OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1 3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18

OpenStudy (mathmale):

One problem per post, please. Which one will it be for this post?

OpenStudy (anonymous):

6x + 2y = 0 4x + y = -1

OpenStudy (mathmale):

You are to use the subst. method. Choose one of these 2 equations and solve it for y. Which equation would you choose for this, and why?

OpenStudy (anonymous):

6x + 2y = 0 This one seems a bit more simpler

OpenStudy (mathmale):

OK. But I'd choose the 2nd one, because solving for y is easier than solving an equation in 2y for y. so: Let's solve 6x+2y=0 for y. Start by dividing all 3 terms by what number?

OpenStudy (anonymous):

6? That or zero.

OpenStudy (mathmale):

Can't divide by zero, so zero is out. I'd suggested that we solve the 1st equation for y. That means focusing on the y term, not the x term. Divide all 3 coefficients of your equation by 2. What do you obtain?

OpenStudy (anonymous):

3x+y=0

OpenStudy (mathmale):

Good. Now, please solve that resulting equation for y.

OpenStudy (mathmale):

All you have to do is to subtract 3x from both sides of the equation.

OpenStudy (anonymous):

So subtract 3x from 3x and 0?

OpenStudy (mathmale):

Yes. 3x-3x=0;. 0-3x = -3x

OpenStudy (mathmale):

What's left?

OpenStudy (anonymous):

y=-3?

OpenStudy (mathmale):

You have to include that ' x ' also. 3x + y = 0 -3x -3x --------------- y= ?

OpenStudy (anonymous):

-3x

OpenStudy (mathmale):

Better. y=-3x. what is the second equation? Type it out now, as is.

OpenStudy (anonymous):

4x + y = -1

OpenStudy (anonymous):

Oh wait, sorry, let me fix it

OpenStudy (anonymous):

4x + -3x = -1

OpenStudy (mathmale):

it's best not to write " +- " Combine these 2 signs: (-1)(1)=?

OpenStudy (mathmale):

Then re-write your equation.

OpenStudy (anonymous):

Im a bit confused by what your asking, really sorry

OpenStudy (mathmale):

You have written 4x + -3y = -1. The " + - " is not acceptable; simplify it to just - .

OpenStudy (anonymous):

Oh ok. 4x -3x = -1 like that?

OpenStudy (mathmale):

Much better. Now combine the 2 x terms on the left side of your equation.

OpenStudy (anonymous):

-1x=-1

OpenStudy (mathmale):

But isn't 4x - 3x = +x ?

OpenStudy (anonymous):

so.... -1x=x? Im sorry this is taking so long, I got really nobody around to explain any of this to me

OpenStudy (mathmale):

that's what I'm here for. ;) 4x - 3x = +x. Therefore your equation 4x - 3x = -1 becomes x = -1. OK with this?

OpenStudy (anonymous):

Im pretty sure.

OpenStudy (mathmale):

We've already found that y=-3x. To find the numerical value of y, please substitute (-1) for x in this equation and compute the value of y. y=?

OpenStudy (anonymous):

-4 - 3 = +x.

OpenStudy (anonymous):

Srry if thats wrong

OpenStudy (mathmale):

formula for y: y=-3x. Known value for x: x=-1 substituting x=-1 into y=-3(x), y=-3(-1) = ? (Note: Multiply, not + or - )

OpenStudy (anonymous):

Im getting confused :(

OpenStudy (mathmale):

I'm sorry. We're going to go thru this problem again. First, let's do it a different way, that might be faster.

OpenStudy (anonymous):

Ok

OpenStudy (mathmale):

6x + 2y = 0 is the first equation; 4x + y = -1 is the second.

OpenStudy (mathmale):

If we use the second equation instead of the first, we won't have to divde the y term by 2. The first equation has a y-term of 2y; the second equation has a y-term of just y. So, it makes sense to solve the 2nd equation for y, not the first.

OpenStudy (anonymous):

Ok

OpenStudy (mathmale):

If the 2nd equation is 4x + y = -1, all we have to do is to subtract 4x from both sides. Can you do that? 4x + y = -1 -4x = -4x ------------------- Please combine these equations to obtain a simpler, new one.

OpenStudy (anonymous):

x+y=3x? or do we just remove that first x?

OpenStudy (anonymous):

Oh wait no.

OpenStudy (anonymous):

x+y=-5x

OpenStudy (mathmale):

Note: 4x -4x ------ 0. Therefore, we get 4x + y = -1 -4x -4x ----------------- 0 = y = -1-4x or y= -1 -4x.

OpenStudy (mathmale):

are you OK with this, or does it need further discussion?

OpenStudy (anonymous):

So right now we got 0 - y = -5x?

OpenStudy (mathmale):

that's 0 + y = -1 - 4x, or simply y = -1 - 4x.

OpenStudy (mathmale):

The -1 and the -4x are NOT like terms, so can't be combined. -1 - 4 = -5 YES but -1 -4x is not -5x; it stays as -1 - 4x.

OpenStudy (anonymous):

Oh ok. And we already know what x is from the other problem

OpenStudy (mathmale):

Don't remember! We have to substitute this y = -1 -4x into the other equation, so we end up with 6x + 2 (-1-4x) = 0 which is an equation in x alone (no y). 6x - 1 - 4x = 0. OK? If not, ask questions.

OpenStudy (anonymous):

Can you explain that in a bit easier way?

OpenStudy (mathmale):

Sorry...I've made a number of typing mistakes which probably have made this even more difficult for you. I'd suggest we just stick with our previous solution and move on to another problem.

OpenStudy (anonymous):

Oh ok. So for this problem can you paste into a reply what I put down? all these different ones are confusing me :(

OpenStudy (mathmale):

We got 4x - 3x = -1, which translates into x = -1. Then, knowing that y=-3x, we got y =3. so, the solution is (-1,3). I'd love to copy and paste, but my computer is not allowing me to do so at the moment. I will go thru this problem with yo u again later if y ou like. but for now, would you please pick anotehr problem to work on? How many such problems do you have left to do?

OpenStudy (anonymous):

3x + 2y = -4 x + 2y = -8 After this one I got 3 more

OpenStudy (mathmale):

Look at the 2nd equation. Since the coefficient of x is 1, I think we should solve this equation for x, not the other equation. if x + 2y = -8, let's subtract 2y from both sides to isolate the y term: -2y -2y -------------- x = ??

OpenStudy (anonymous):

-6y?

OpenStudy (mathmale):

We have -8 - 2y on the right side. -8 and -2y are NOT "like" terms in that one has y, the other does not. Thus, we cannot combine the -8 and the 2 to get 6. Again, the final result is x= -8 - 2y. Can you accept this? If not, ask questions.

OpenStudy (mathmale):

@Pingurox ?

OpenStudy (anonymous):

Sorry, had to feed my dog

OpenStudy (anonymous):

And yes, I can accept that

OpenStudy (mathmale):

Since you can review this conversation later, please accept that x= -8 -2y. Next, we substitute this into the other equation: 3x + 2y = -4.

OpenStudy (mathmale):

3( ) + 2y = -4 Please write -8-2y inside this set of parentheses.

OpenStudy (anonymous):

3(-8-2y ) + 2y = -4

OpenStudy (mathmale):

Yes. Good. Multiply each term -8 -2y by 3. What do you get? (Leave the other terms alone)

OpenStudy (anonymous):

So multiply -8 -2y by 3 each?

OpenStudy (mathmale):

Yes. that's the "Distribution Rule of Multiplication."

OpenStudy (anonymous):

So I multiple them seperately or together?

OpenStudy (mathmale):

Multiply the -8 by 3. Separately, multiply the -2y term by 3.

OpenStudy (mathmale):

Separately. And keep the products (answers) separate.

OpenStudy (anonymous):

-8 by 3.= -24 -2y by 3. =-6y

OpenStudy (mathmale):

Right. So, on the left side of your equation, you have -24 - 6y +2y = -4. Combine the -6y and the -2y on the left, to form just one term with y in it.

OpenStudy (mathmale):

-6y+2y = ? (note that you have 2 different signs here).

OpenStudy (anonymous):

Uhh....im a wee bit lossed

OpenStudy (mathmale):

-6+2=-4, so -6y+2y = -4 y. Then, on the left side y ou have -24 - 4y=-4. Adding 24 to both sides eliminates the -24 on the left and results in -4y=-4+24, or -4y=20 Can y ou now find y?

OpenStudy (mathmale):

If -4y=20, divide both sides by -4.

OpenStudy (anonymous):

divide -4y and 20 by -4?

OpenStudy (mathmale):

Yes. -4y divided by -4 is simply y. We want y alone on the left side. So, now on the left side we have y, and on the right side we have 20 / (-4), or

OpenStudy (mathmale):

\[\frac{ 20 }{ -4 }=?\]

OpenStudy (mathmale):

What is 20 divided by -4?

OpenStudy (mathmale):

20 divided by 4 is 5. The only thing different in our "what is 20 divided by -4" is that the 4 is negative. 20/4=5, 20 / (-4) = ??

OpenStudy (anonymous):

-5

OpenStudy (mathmale):

right. So, on the left of our equation we now have y and on the right side we have -5. Therefore, y = -5. You can substitute y=-5 into either of the 2 given equations to find x. suppose we choose to use the equation x+2y=-8. Substituting y = -5, x + 2(-5) = - 8 What is 2(-5), if we multiply?

OpenStudy (anonymous):

-10

OpenStudy (mathmale):

Yes, and so you have x - 10 = -8 and you need to add 10 to both sides to obtain a value for x. Finish this, please: x - 10 = -8 +10 = +10 ________________ x = ?

OpenStudy (anonymous):

x-20=-18 Im just gueesng....x=-2?

OpenStudy (mathmale):

On the right side you have -8 + 10, and that comes out to just 2. Therefore, x = 2. Earlier, we found that y =-5 Therefore your solution is (2,-5). x=2 and y=-5. These 2 equations have graphs which are straight lines. If you were to graph them you'd see that they intersect at the point (2,-5).

OpenStudy (anonymous):

Alright. So just copy paste that into my worksheet for my solution?

OpenStudy (mathmale):

You can do that, of course. But I'd feel better if you'd not just accept this result as our answer, but rather check it by substitution. You can substitute x=2 and y=-5 into either equation. If the equation is then true, your solution is correct; otherwise it is not correct.

OpenStudy (mathmale):

I encourage you to try this! I've done it myself and find that the first equation is true when x=2 and y=-5. But it's important that y ou know how to do this for ourself.

OpenStudy (mathmale):

@Pingurox: I'm sorry but my head is swimming after hours on OpenStudy. I need to get off OpenStudy. I hope you'll go through our solution of the 2nd problem again and make a list of any questions you may have about that process. We could discuss your questions later. But I have to excuse myself right now. Best of luck to you, and thanks for your persistence.

OpenStudy (anonymous):

No problem. I gtg anyways

OpenStudy (anonymous):

anybody still there